OpenStudy (anonymous):

help D: What is the constant of variation for the quadratic variation? 6y = 4x^2

Nnesha (nnesha):

y =k x^2 quadratic equation for direct variation wher k is constant of variation

Nnesha (nnesha):

where*

Nnesha (nnesha):

so solve for y :)

OpenStudy (anonymous):

6y = 4x^2 y = (4/6)x^2 y = (2/3)x^2

OpenStudy (anonymous):

Since y is being multiplied by 6 and you want y by itself

OpenStudy (anonymous):

So divide both sides by 6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):